OpenStudy (anonymous):

Need help solving this please 2x^2 + 11x + 2 = -13

OpenStudy (anonymous):

I think I have to add 13 to each side

OpenStudy (anonymous):

so it can = 0?

OpenStudy (superdavesuper):

sure that is the right first step. keep going :)

OpenStudy (anonymous):

so 2x^2 + 11x + 15 = 0 I don't really know what to do from here

OpenStudy (superdavesuper):

U know how to solve a quadratic eqn: ax^2 + bx + c = 0?

OpenStudy (anonymous):

Not really. I think I'm supposed to factor and get two separate equations

OpenStudy (anonymous):

but not really sure what steps to take

OpenStudy (superdavesuper):

Well if u dont have the general solution, then yes u will have to try to factor them in....

OpenStudy (superdavesuper):

2x^2 = 2x * x 15 = 15 * 1 or 5 * 3 So u need to figure out how to factor them so u can get 11x...good luck!